### v1

# GraphMend: Code Transformations for Fixing Graph Breaks in PyTorch 2

Abstract
This paper presents GraphMend, a high-level compiler that eliminates FX graph breaks in PyTorch 2 programs. Although PyTorch 2 introduced TorchDynamo and TorchInductor to enable just-in-time graph compilation, unresolved dynamic control flow and unsupported Python constructs often fragment models into multiple FX graphs. These fragments force frequent fallbacks to eager mode, incur costly CPU-to-GPU synchronizations, and reduce optimization opportunities. GraphMend addresses this limitation by analyzing and transforming source code before execution. Built on the Jac compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python I/O functions. This design allows PyTorch’s compilation pipeline to capture larger, uninterrupted FX graphs without requiring manual refactoring by developers. Evaluation across eight Hugging Face models shows that GraphMend removes all fixable graph breaks due to dynamic control flow and Python I/O functions, driving the break count to 0 in 6 models and reducing it from 5 to 2 in another model. On NVIDIA RTX 3090 and A40 GPUs, GraphMend achieves up to 75% latency reductions and up to 8% higher end-to-end throughput. These results demonstrate that high-level code transformation is an effective complement to PyTorch’s dynamic JIT compilation pipeline, substantially improving both usability and performance.
I Introduction
PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility [1]. However, its general-purpose design incurs efficiency overheads, leaving PyTorch workloads less optimized than handcrafted GPU implementations. To bridge this gap, PyTorch 2 [2] introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
One of the central intermediate representations (IR) in this pipeline is the FX graph [3]. TorchDynamo intercepts and symbolically evaluates Python bytecode to create an FX graph. The FX graph is then passed as the IR to a backend compiler, by default TorchInductor [2], which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
A key component in the PyTorch compilation pipeline is the capture of code into an FX graph through symbolic evaluation. It is important to avoid splitting the forward function into smaller graphs, since unsupported code triggers execution to fall back to eager mode, fragmenting the computation. For example, TorchDynamo often encounters difficulties when resolving conditional jumps, as control flow depending on dynamic values cannot be statically determined [4]. In such cases like dynamic control flow or calls to certain Python built-ins, TorchDynamo cannot capture the affected code into the FX graph. Instead, it inserts a graph break, which splits the computation into multiple fragments [2]. As a result, a single forward function may be compiled into several disjoint FX graphs rather than one unified graph.
The compiled FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode [1]. Each fallback to eager execution requires GPU kernels to synchronize with the CPU to fetch and schedule subsequent operations. These repeated host-device synchronizations introduce substantial overhead, since control is repeatedly transferred between CPU and GPU instead of executing long fused kernels entirely on the device. Moreover, graph breaks reduce optimization opportunities, as each FX graph is compiled in isolation, preventing cross-graph fusion and global optimization [5, 2].
Avoiding graph breaks is essential for realizing the full performance potential of PyTorch 2. The current PyTorch compilation pipeline begins at the bytecode level and does not exploit source-level information. As a result, when high-level constructs such as conditional branches are lowered into jump instructions at the bytecode level, TorchDynamo often cannot statically resolve them, causing graph breaks. Our key insight is that compilers should not limit transformations to the bytecode level. Instead, more holistic compilation strategies that analyze and transform higher-level program representations closer to the source code can expose additional optimization opportunities. By incorporating source-level analysis and transformation, high-level code regions can be systematically rewritten into forms that enable more effective optimization and unified graph generation within the PyTorch 2 pipeline.
Building on this insight, we introduce GraphMend, a high-level compiler technique for PyTorch 2 models that eliminates two common types of graph breaks those caused by dynamic control flow and Python I/O operations through program analysis and transformation. Unlike standard PyTorch 2, which captures execution at runtime to construct graphs for JIT compilation, GraphMend restructures Python code before execution. We implement GraphMend within the Jac framework [6, 7, 8], a language infrastructure that provides compiler and runtime support for Python programs. The Jac compiler constructs an abstract syntax tree (AST) to capture program structure and a control-flow graph (CFG) to represent execution paths, merging them into a unified IR [9]. Multiple compiler passes operate on this IR to analyze and transform the program, and GraphMend extends this pipeline with additional passes that detect code patterns likely to cause graph breaks. Specifically, GraphMend introduces compiler passes that apply two key transformations: Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects, which systematically eliminate graph breaks. The transformed program is then compiled into standard Python bytecode and executed by the regular CPython interpreter, seamlessly integrating with the existing PyTorch 2 compilation pipeline. This design enables GraphMend to automatically restructure code before it reaches TorchDynamo.
We evaluate GraphMend on a benchmark suite of 8 Hugging Face [10] models that trigger FX graph breaks. GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed. On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput. Improvements are larger when a graph break lands in the middle of a hot function where its disruption is highest, and smaller when the break is less disruptive; very small models also tend to see larger relative reductions in per forward pass latency because CPU to GPU handoff dominates a greater share of their runtime.
Our contributions are, (1) We introduce a benchmark suite of eight Hugging Face models that naturally exhibit graph breaks, providing a realistic and reproducible testbed for evaluating PyTorch compilation. (2) We design and implement the GraphMend compilation technique, which automatically restructure high level code to eliminate common graph breaks. (3) We conduct comprehensive evaluations on this benchmark suite using real world models and hardware, demonstrating the effectiveness of our approach in reducing graph breaks and improving performance.
II Background
Optimization pipelines in PyTorch aim to capture execution graphs that can be globally optimized by compilers, unlike eager mode where each operation runs immediately and independently [1, 3].
II-A Graph Capturing Techniques in PyTorch Before PyTorch 2
Early approaches to graph capture were limited. torch.jit.trace [11] used record–replay at the C++ dispatcher level but ignored Python control flow and often produced incomplete or misleading graphs. torch.jit.script [12] analyzed Python ASTs but followed an “all-or-nothing” rule: if a single construct was unsupported, capture failed entirely. To address these limitations, PyTorch introduced the torch.fx [3] module, enabling symbolic execution of Python code to generate an FX graph, a higher-level IR. Although more flexible, symbolic_trace could still misrepresent certain operations.
II-B FX Graph IR
An FX graph represents operations as nodes connected by data dependencies [2, 3]. Unlike eager execution, which evaluates operations sequentially, an FX graph exposes the computation as a data flow structure. This global view enables the compiler backend to apply optimizations such as operation fusion and kernel specialization.
To illustrate, consider the simple PyTorch forward function in Figure 1(a). When captured by TorchDynamo, it is transformed into the FX graph shown in Figure 1(b), where each operation is represented as a node in the IR.
1def forward(x, y):
2 z = x*2 + y*2
3 return torch.relu(z)
(a) Forward function in PyTorch
%x : torch.Tensor
%y : torch.Tensor
%mul_x : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%x, 2)
%mul_y : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%y, 2)
%z : torch.Tensor = torch.ops.aten.add.Tensor(%mul_x, %mul_y)
%out : torch.Tensor = torch.ops.aten.relu.default(%z)
return (%out)
(b) Corresponding FX graph IR
Figure 1: A simple forward function (a) and its compiled FX graph IR (b).
This FX graph serves as the foundation for subsequent compiler passes that optimize and lower the computation into efficient GPU kernels.
II-C TorchDynamo in PyTorch 2
PyTorch 2 [2] introduced TorchDynamo, Python bytecode-level tracer tightly integrated with CPython. It symbolically evaluates Python bytecode to extract PyTorch operations into FX graphs, while delegating unsupported constructs back to Python. This avoids incomplete graphs but may still fragment the computation by generating partial graphs by inserting graph-breaks.
1def f(x, y):
2 x_1 = x*2
3 y_1 = y*2
4 if x.sum() > 10: # <-- graph-break here
5 z = x_1 + y_1
6 else:
7 z = x_1 * y_1
8 return torch.relu(z)
(a) A PyTorch forward pass with data dependent control flow. The conditional branch on x.sum cannot be captured, causing a graph break.
1def f(x, y):
2 x_1 = x*2
3 y_1 = y*2
4 cond = x.sum > 10
5 z_add = x_1 + y_1
6 z_mul = x_1 * y_1
7
8 # rewrite dynamic control flow using tensor select
9 z = torch.where(cond, z_add, z_mul)
10 return torch.relu(z)
(b) Eliminating the graph break by rewriting data dependent control flow with torch.where, keeping the computation in a single FX graph.
Figure 2: Comparison of control flow handling in torch.compile: (a) graph break due to Python control flow, and (b) fixed version using torch.where.
Figure 3: Profiled traces of forward pass execution across CPU and GPU. (a) Forward pass execution trace of code with graph breaks in Figure 2(a). (b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2(b).
III Motivating Example
As discussed in § II, TorchDynamo attempts to capture a single FX graph for the forward function but inserts graph breaks whenever it encounters unsupported constructs, forcing execution to fall back to eager mode [2].
III-A FX graph breaks
Let us consider the modified program in Figure 2(a), which adds a conditional branch to the forward pass of Figure 1(a). When Dynamo tries to symbolically evaluate the function to capture FX graph, it cannot capture the conditional branch because the outcome depends on the runtime value of the tensor expression x.sum. Therefore, it inserts a graph break. As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph [2, 13].
III-B Impact of graph breaks
The backend compiler (TorchInductor by default) compiles each captured FX graph separately and lowers it to GPU executables such as fused Triton kernels [2, 14]. Each compiled FX graph corresponds to a torch.compiled region. When graph breaks occur, a single Python function results in multiple compiled regions separated by segments of unsupported Python code [13].
1def fn(x):
2 x = torch.relu(x)
3 print("tensor:", x) # <-- graph-break here
4 return torch.sin(x)
(a) Compiled function with a debug print of tensor statistics. Dynamo inserts a graph break at the print statement.
1def fn(x):
2 x = torch.relu(x)
3 to_print = "tensor:", x # <-- assign variable to print
4 y = torch.sin(x)
5 print(to_print) # <-- print at the end of tracing
6 return y
(b) Fixing the graph break by registering print as a reorderable logging function. TorchDynamo can now safely reorder the call and keep the computation in a single FX graph.
Figure 4: Fixing graph breaks due to Python I/O: (a) direct print causes a graph break, (b) reordering via variable assignment avoids the break.
1 if x > 0:
2 return 1
3 else:
4 return -1
(a) Python function with if-else branch
2 LOAD_CONST 1 (0)
4 COMPARE_OP 4 (>)
6 POP_JUMP_IF_FALSE 14
8 LOAD_CONST 2 (1)
10 RETURN_VALUE
14 LOAD_CONST 3 (-1)
16 RETURN_VALUE
(b) Disassembled Python bytecode
1 if x > 0:
2 return 1
3 return -1
(c) Python function having identical bytecode to (a)
Figure 5: Bytecode-level limitations of TorchDynamo.(a) (a) and (c) show two syntactically different functions with equivalent semantics. (b) Both compile to identical Python bytecode.
During runtime, the first compiled region runs on the GPU, after which control returns to the Python interpreter to execute unsupported code in eager mode before switching back to the GPU for subsequent compiled regions. These context switches introduce overhead and lower GPU utilization per forward pass [2].
This effect is illustrated in Figure 3(a), which shows the CPU–GPU profiling trace of the function in Figure 2(a). The trace reveals two separate torch.compile regions, corresponding to two FX graphs. The graph break is visible in the device-to-host (D2H) memcpy following the first FX graph, where control returns to the CPU to execute unsupported code, leaving the GPU idle until the second CUDA graph launch begins. This fragmentation both increases CPU-side scheduling overhead and prevents cross-graph optimizations such as kernel fusion.
If the graph break is eliminated and the computation is captured in a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA graph region. In this case, the GPU remains fully utilized without falling back to the CPU, and no device-to-host memory transfers are observed between regions. As a result, kernel launches are streamlined and synchronization overhead is removed. Therefore, eliminating graph breaks is necessary.
IV GraphMend Compiler Technique
Figure 6: GraphMend compiler integration in the Jac pipeline. The pipeline (top) lowers Python/Jac source code into a unified intermediate representation (UniiR) with AST, symbol table, and CFG. GraphMend (bottom) analyzes entry points, detects graph breaks, and applies AST transformations (Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects) to eliminate breaks before execution.
As discussed in § III, graph breaks in PyTorch 2 introduce significant overhead, undermining the performance gains expected in PyTorch 2. In this work, we focus on two common sources of graph breaks and introduce a compiler technique that eliminates them prior to execution.
IV-A Classification of Graph Breaks
In practice, most graph breaks arise from two common situations:
  1. 1.
    Data dependent operations [2, 13]: Constructs whose outcome depends on runtime values and hence, cannot be traced.
  2. 2.
    Unsupported operations [2, 13]: Several Python built in functions or certain functionalities that TorchDynamo does not support.
In this section, we discuss these types of graph breaks.
IV-A1 Data dependent operations
Dynamo inserts graph breaks when it encounters control flow that depends on tensor values, such as if-statements or loops, as well as direct tensor data accesses (e.g., .item, .data_ptr). Among these cases, when Dynamo encounters dynamic control flow based on a tensor value as shown in Figure 2(a), we can fix the graph break by rewriting the code into operations supported by the GPU.
IV-A2 Python builtin functions
Some Python built-in functions such as printing, logging, or issuing warnings can also trigger graph breaks. Figure 4(a) shows an example where a debug print statement in the middle of the function causes a break.
IV-B Fixing Common Graph Breaks
Among these graph break reasons, there are two main reasons that are common and fixable which are (1) Data-dependent Control Flow and (2) Python I/O Operations.
IV-B1 Data-dependent Control Flow
Dynamo inserts graph breaks for constructs such as if-statements or loops where branching depend on tensor values. Returning to the example in Figure 2(a), the conditional introduces a graph break. In such cases,we can rewrite the function using tensor operations such as torch.where, which are supported by the GPU. The graph break fixed version using torch.where is shown in Figure 2(b) for the same function in Figure 2(a). This eliminates the graph break and allows Dynamo to capture a single FX graph.
IV-B2 Python I/O Operations
Python I/O functions such as Printing/logging/issuing warnings will result in a graph break. An example for this is shown in Figure 4(a) as there is a debug print statement in the middle of the function. If we can defer print statement to the end of the function, then we can avoid graph breaks as shown in Figure 4(b).
These graph-fixing techniques are simple and straightforward: by applying source-level transformations, we can often avoid graph breaks. However, TorchDynamo traces at the bytecode level, which prevents it from leveraging such transformations and leads to unnecessary graph breaks. For example, consider Figure 5. The function in Figure 5(a) uses an explicit if–else block, while the function in Figure 5(c) omits the else and relies on a fall-through return. Although the two versions differ syntactically, both compile to the same bytecode shown in Figure 5(b). Because TorchDynamo operates purely on bytecode, it cannot distinguish between these semantically equivalent but syntactically distinct patterns.
Similarly, logging operations introduce host I/O into the forward pass. Because it is unsafe to alter such effects directly at the bytecode level, TorchDynamo also treats these sites as graph breaks. In both cases, the limitation stems from relying solely on bytecode-level semantics captured just-in-time.
To address this, we implement a compilation technique that analyzes and transforms programs at a higher-level IR. By operating on high-level IR, we preserve structural information lost in bytecode, enabling more sophisticated analyses and safe source-to-source rewrites that remove graph breaks before execution. This implementation of the technique is shown in Figure 6 which is elaborated in the rest of this section.
IV-C Jac Compilation Pipeline
To implement our compiler technique, we build on the Jac framework [6, 7, 9], which extends Python with additional language features through its runtime system. The Jac compiler accepts both Jac and Python code, parses it into an abstract syntax tree (AST), and progressively lowers it into a unified intermediate representation (UniiR). As illustrated in the top half of Figure 6, UniiR interleaves three structures: AST, control-flow graph (CFG), and symbol table. In this way, structural, semantic, and control-flow information are combined in a single representation. This integration enables cross-cutting analyses that are difficult to achieve when these structures are maintained separately. The CFG is constructed by linking AST nodes to their successors, embedding control-flow edges within the tree. This process is implemented in the CFGBuildPass.
Our GraphMend pass takes UniiR as input and performs analyses to identify the two types of graph breaks, as illustrated in the bottom half of Figure 6. It then applies transformations on the AST to rewrite the code and eliminate these breaks. Afterward, UniiR is rebuilt and verified before being passed back to the Jac compiler pipeline. At that stage, the pipeline can either compile the program to bytecode for execution on the Python interpreter or generate a transformed Python source file.
Input: UniiR; TorchAttrTable (torch attributes $\mapsto$ {dynamic, static})
Output: Tags: GraphBreak[DynCtrlFl], GraphBreak[logger/print]
foreach Dynamo Entry Point $r$ in UniiR do
  Initialize worklist with successors of $r$
  while worklist not empty do
     Pop node $n$ from worklist; push unseen successors of $n$
    // 1) Dynamic control-flow via if-conditions
     if $n$ is an IfStmt then
        $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
        $dyn\leftarrow\textbf{false}$
        foreach $e\in E$ do
           if $e$ has a torch attribute then
              $a\leftarrow$ attribute of $e$
              // e.g., .sum
              if TorchAttrTable[$a$] = dynamic then
                 $dyn\leftarrow\textbf{true}$; break
              end if
           end if
        end foreach
       if not dyndyn then
           if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
              $dyn\leftarrow\textbf{true}$
           end if
        end if
       if dyndyn then
           Tag $n$ as GraphBreak[DynCtrlFl]
        end if
     end if
    // 2) Logger / print side effects
     if $n$ is a call to print or logger.* then
        Tag $n$ as GraphBreak[logger/print]
     end if
  end while
end foreach
Algorithm 1 Fixable Graph-Break Detection
IV-D GraphMend Implementation
Our compiler technique consists of three main passes, illustrated in the bottom half of Figure 6.
IV-D1 Dynamo Entry Point Analysis Pass
To identify potential TorchDynamo compilation entry points, this pass traverses the AST and analyzes each node. A node is tagged when a function or an nn.Module object wrapped or decorated with torch.compile is encountered.
If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped. Otherwise, the tagged nodes serve as entry points for subsequent passes.
IV-D2 Graph Break Type Analysis Pass
We implement a compound analysis pass on UniiR to identify potential graph-break locations, following the procedure in Algorithm 1. Starting from each Dynamo entry point, the pass traverses the control-flow graph (CFG) to locate program points that may induce graph breaks. For conditional branches, the pass inspects the subexpressions of the condition. If any involve Torch attributes known to be dynamic (e.g., .sum()), the branch is marked as dynamic control flow. Otherwise, a use-def analysis over the symbol table and CFG determines whether the condition depends on model inputs or tensor values; if so, it is also tagged as dynamic. In addition, side-effecting operations such as print and logger.* calls are detected and annotated as logging-induced graph breaks as in Algorithm 1.
These annotations guide the subsequent transformation passes that eliminate graph breaks.
IV-D3 AST Transformation Pass
After identifying graph-break locations, the AST must be modified to eliminate them. As discussed earlier in this section, these fixes can be applied through source-level transformations. Following the same approach, the AST is modified according to the graph-break tags identified in the previous pass. Based on these tags, one of the following transformations is performed.
Predicated Dynamic Control Flow Transformation
To address data-dependent control-flow graph breaks, we first evaluate the branch condition and store it as a predicate. The control flow is then rewritten using torch.where, which enables predicated execution of both branches. This process is illustrated in Figure 6. In the AST Transformation Pass (Figure 6), the original branch if x.sum() > 5: ... else: ... is reduced by extracting the condition into a symbolic variable pred (R 1). Next, the two branch assignments are merged into a single dataflow expression using torch.where(pred, e1, e2). Finally, the if and else AST nodes are removed, and their child nodes are reconnected to the parent, producing a transformed tree in which both execution paths are explicitly represented.
Graph-Epilogue Deferred Side Effect Transformation
For print or logger calls, the message is first stored in a temporary variable. This variable is then passed to the outermost forward function and executed at the end of the function (epilogue) before returning, as shown in Figure 6. In the example ASTs in Figure 6, the original call print("Logger message") is rewritten as a deferred assignment deferred = "Logger message". This assignment is then moved to the epilogue block, where it is printed after graph execution (R 2). In this way, the logging effect is preserved while remaining safely outside the traced region. Note that any computation in the return statement must be moved before the epilogue, stored in a temporary variable, and then returned afterward to avoid introducing a graph break.
IV-E Healing UniiR and Downstream Flow
After the AST transformations, the UniiR is no longer consistent. Therefore, the symbol table and CFG must be rebuilt using the same passes that originally created them. As shown in Figure 6, once the GraphMend pass is complete, the SymTabBuildPass and CFGBuildPass are rerun to restore UniiR for downstream compilation.
At the end of the Jac compiler pipeline, Python bytecode is generated for interpretation, enabling TorchDynamo to seamlessly operate on the transformed code.
TABLE I: Models in benchmark suite
Benchmark Model Name Graph break count Graph break reasons Total Operators
biogpt [15] 2 logger calls 408
blenderbot-400M-distill [16] 3 logger calls 489
flan-t5-large [17] 3 logger calls 2130
longformer-base-409 [18] 5 logger calls, tensor.item() call 913
moe-minicpm-x4-base [19] 15 Dynamic shape operator 280
Phi-4-mini-instruct [20] 5 Dynamic control flow 1541
Qwen-Audio-Chat [21] 2 Dynamic control flow 1647
tiny-random-PegasusForCausalLM [22] 2 logger calls 64
V Evaluation
We evaluate GraphMend on real-world models that exhibit graph breaks. Our goals are to quantify (i) how often GraphMend removes fixable breaks and (ii) the end-to-end improvement after those fixes.
Experimental Setup
We implement GraphMend on top of Jaseci [6, 7, 9], an open-source research infrastructure that provides both compiler and runtime control as a Python superset. Our experiments are conducted on NVIDIA RTX 3090 and A40 GPUs, using TorchInductor as the backend compiler.
Benchmark Suite
To construct our benchmark, we (1) randomly sample 65 Hugging Face [10] models spanning diverse architectures and task categories; (2) identify models that trigger FX graph breaks; and (3) retain the eight models with breaks for evaluation. Table I summarizes the resulting suite, including the number and causes of graph breaks and the total operator count. A similar methodology was used in prior works to construct benchmark suite [2].
Profiling Methodology
We profile the forward pass using PyTorch Profiler [23]. For each model, we evaluate two configurations: original model and graph breaks fixed version. Each configuration is executed for 7 iterations: one cold-start iteration followed by six warm iterations. The profiler records CPU and CUDA activities and operator-level events; traces are exported and inspected with Chrome’s trace viewer.
V-A GraphMend’s Ability to Fix Graph Breaks
TABLE II: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
Benchmark Model Graph Breaks Fixed (%)
biogpt 2 100
blenderbot-400M-distill 3 100
flan-t5-large 3 100
longformer-base-4096 5 40
moe-minicpm-x4-base 15 0
Phi-4-mini-instruct 5 100
Qwen-Audio-Chat 2 100
tiny-random-PegasusForCausalLM 2 100
First, we check GraphMend’s ability to fix graph breaks across the benchmark suite. Table II shows that GraphMend fully eliminates graph breaks in 6 of the 8 models. It partially fixes longformer-base-4096 and leaves moe-minicpm-x4-base unchanged. By cause, GraphMend removes all breaks due to logging statements and data-dependent control flow. The remaining breaks: tensor.item() and dynamic-shape operators could not be fixed by code transformations. For the performance analysis that follows, we report results on the seven models whose break counts were fully eliminated or reduced, excluding moe-minicpm-x4-base.
V-B Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
We measure Relative Latency Improvement (%) as
$100\times\frac{T_{\text{break}}-T_{\text{fixed}}}{T_{\text{break}}},$
where $T_{\text{break}}$ is the per-forward-pass latency for the benchmarked model with graph breaks, and $T_{\text{fixed}}$ is the latency after applying GraphMend fixes.
We report the relative improvement per forward pass for break-fixable models in the benchmark suite, using PyTorch Profiler data for both cold and warm (steady-state) runs.
Cold Start Latency Improvement
Figure 7(a) reports results on both RTX 3090 and A40 GPUs. Across models, we observe a relative latency reduction of about 30–75% for both GPUs, indicating substantial gains in cold runs. At cold start, each graph break incurs significant overhead because every CUDA graph requires a separate record and caching. After eliminating these breaks, only a single (or fewer) CUDA Graph capture is needed, resulting in fewer records and reduced capture time, which minimizes overhead. Consequently, we achieve larger relative latency improvements per forward pass in cold runs. A more detailed analysis is provided in  § V-D.
Steady-State Latency Improvement
Figure 7(b) shows steady-state improvements on RTX 3090 and A40 GPUs, ranging from about 2.5% to 25% across models. After the cold run, the captured graph is cached and replayed, removing the high capture overhead present initially. However, when graph breaks persist, execution still incurs GPU–CPU switches (as discussed in  § V-D), which add latency. Eliminating these breaks with GraphMend removes such switches, yielding additional improvements in relative latency per forward pass.
Notably, tiny-random-PegasusForCausalLM shows over 25% improvement. Since this model is very small, with only 64 operators, the fixed costs from graph breaks (such as host–device synchronization, and extra graph launches) take up a large portion of the runtime. Removing these breaks therefore gives a huge speedup per forward pass.
Figure 7: Latency improvements from GraphMend (a) Cold-start latency reductions (b) Steady-state latency reductions across benchmark models on RTX 3090 and A40 GPUs.
Figure 8: Relative throughput improvement across benchmark suite
V-C Relative Throughput Improvement by Fixing Graph Breaks
We measure Relative Throughput Improvement (%) as
$100\times\frac{Tp_{\text{fixed}}-Tp_{\text{break}}}{Tp_{\text{break}}},$
where $Tp_{\text{break}}$ is the throughput (output tokens per second) for the model with graph breaks, and $Tp_{\text{fixed}}$ is the throughput after applying GraphMend fixes.
We measure the relative throughput across the models in the benchmark suite that are fixable through GraphMend for both RTX 3090 and A40. Figure 7(b) shows a relative throughput improvement of about 5-8% across all models on both GPUs. This improvement follows from the relative latency reduction per forward pass observed in Section V-B: as each step becomes faster, the end-to-end tokens per second increase. The variation of improvement across models highlights an important point: the magnitude of improvement depends on the extent to which a graph break affects the forward path. When the impact of a break is small, the corresponding throughput gain is also limited.
For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow. We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct from both GPUs. The difference arises because the break in Phi-4-mini-instruct occurs near the beginning of the forward function, whereas in Qwen-Audio-Chat it occurs in the middle, where its impact is higher. Thus, the location of a break determines its effect: fixing some breaks yields larger throughput gains, while fixing others has a smaller impact.
tiny-random-PegasusForCausalLM achieves more than 25% lower forward-pass latency, but throughput gains are smaller. This is because throughput also depends on prefill, decode, CPU tasks (e.g., sampling, tokenization), and data transfers, which are unaffected. For such a small model, these fixed costs dominate, so removing the break speeds up GPU compute but has less impact on end-to-end tokens/s, especially when decode dominates generation.
V-D Overhead Analysis of Graph Breaks
We conducted an analysis to evaluate how fixing graph breaks with GraphMend reduces overhead. We analyze Qwen-Audio-Chat model running on an A40 GPU which has 2 graph breaks.
(a) Profiler tracer for one iteration of warm runs for Qwen-Audio-Chat original model before fixing the graph break run in A40 GPU
(b) Profiler tracer for one iteration of warm runs for Qwen-Audio-Chat model after fixing the graph break run in A40 GPU
Figure 9: (a) Profile traces.
Figure 10: Profiler tracer of cold run of Qwen-Audio-Chat model run on A40 GPU.
(a) Steady State - Original Model
(b) Cold Start - Original Model
(c) Steady State - Fixed Model
(d) Cold Start - Fixed Model
Figure 11: CPU/GPU activity time traces for Qwen-Audio-Chat model run on A40 GPU.
To ensure a consistent comparison in this section, all activity plots (Figures 11(a), 11(b), 11(c), and 11(d)) are drawn only for the time interval between the first and second CUDA graph executions in the original model; for the fixed model, we consider the identical time region.
Steady-state run overhead analysis
Figure 9(a) shows the profiler trace for the original model. Execution is split into three CUDA graphs because of two graph breaks, with idle GPU periods between the first and second FX graph executions. This is clearly visible in Figure 11(a), the GPU becomes idle while the CPU handles control flow in eager mode, including a device-to-host (D2H) memcpy triggered by dynamic control flow (graph break reason for this model). This CPU fallback introduces an overhead.
After applying GraphMend, the rewritten code replaces the dynamic branch with a GPU-supported torch.where. The resulting trace (Figure 9(b)) shows one continuous CUDA graph. As confirmed in Figure 11(c), GPU activity remains uninterrupted and CPU fallbacks disappear. Thus, GraphMend eliminates the overhead of CPU–GPU context switching during steady-state runs.
Cold run overhead analysis
Cold runs introduce even larger overheads. In the profiler trace of original model (Figure 10), each graph break triggers a separate CUDA Graph recording and caching, leading to long idle GPU phases and heavy CPU involvement as shown in Figure 11(b). In this interval, we observe that GPU idle time dominates, while the CPU is heavily active, reflecting significant overhead. By contrast, the fixed model (Figure 11(d)) executes with a single capture, avoiding repeated records and reducing idle GPU time. Overall, GraphMend reduces overhead more significantly in cold runs than in steady state, explaining the substantial cold-start latency reductions per forward pass reported in Figure 7(a).
V-E GPU Runtime Kernel Statistics Analysis
We analyze how fixing graph breaks affects GPU kernel counts, ordering, and scheduling, and whether this enables new optimization opportunities. Using the Phi-4-mini-instruct model on an A40 GPU, the profiler reports 404 executed kernels for the original model and 393 for the fixed model with a reduction of 11. This drop indicates more fusion and fewer launches, as removing breaks allows the compiler to capture larger regions and emit fewer, larger kernels.
Figure 12 illustrates this change. In the original model, 3 small kernels appear that are replaced in the fixed model by a single fused kernel. We also observe reordering: one kernel from the original run shifts upward in the fixed run. With one CUDA Graph, the runtime can reorder independent regions and pack launches more effectively, reducing stalls and improving locality.
Graph breaks force separate CUDA Graph captures, blocking fusion across the break and restricting scheduling to smaller windows. Removing them allows larger portions of the forward pass to be captured together, enabling the compiler and runtime to fuse operations, remove redundant launches, and reorder independent work. As a result, kernel gaps shrink and device utilization improves.
In summary, fixing graph breaks with GraphMend not only reduces launch overhead but also exposes new optimization opportunities, underscoring the importance of eliminating such breaks.
Figure 12: Kernel Fusion and Reordering Visualization for Phi-4-mini-instruct model.
VI Related Work
PyTorch has long sought to balance performance with Python’s flexibility. TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs [12, 11]. PyTorch 2 introduced TorchDynamo for bytecode-level capture and TorchInductor for backend compilation [2, 1], yet graph breaks remain common when encountering data-dependent branches or Python I/O. Our work reduces these breaks by rewriting source code before execution, allowing more continuous FX graphs.
General compiler infrastructures such as XLA [24], TVM [25], and MLIR [26] provide IR-based optimization, while Glow [27], TASO [28], and ONNX Runtime [29] extend portability across hardware and frameworks. These systems assume large graphs are already available, whereas we restructure PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
Dynamic models motivate systems like GRAPE [30] and Triton [31, 32], which manage runtime variability through efficient kernel generation. PyTorch’s compiler stack still falls back to eager execution in such cases, while our method removes common break sources by rewriting branches into tensorized forms that remain graph-compatible.
Program transformations more broadly, as seen in Halide [33], Lantern [34], and DLVM [30], show the potential of semantics-preserving rewrites for optimization. Our approach follows this tradition but targets PyTorch specifically, introducing lightweight transformations that enable TorchDynamo to capture larger uninterrupted graphs.
VII Conclusion
We presented GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution. By restructuring code constructs such as dynamic control flow and Python I/O operations, GraphMend reduces fragmentation in the compilation pipeline and enables larger unified graphs to be captured by TorchDynamo. Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable breaks and consistently improves runtime efficiency, with substantial latency and throughput gains on modern GPUs.

### v2

@@
-GraphMend: Code Transformations for Fixing Graph Breaks in PyTorch 2
+
 Abstract
-This paper presents GraphMend, a high-level compiler that eliminates FX graph breaks in PyTorch 2 programs.
+This paper presents GraphMend, a high-level compiler technique that eliminates FX graph breaks in PyTorch 2 programs.
 Although PyTorch 2 introduced TorchDynamo and TorchInductor to enable just-in-time graph compilation, unresolved dynamic control flow and unsupported Python constructs often fragment models into multiple FX graphs.
-These fragments force frequent fallbacks to eager mode, incur costly CPU-to-GPU synchronizations, and reduce optimization opportunities.
+These fragments force frequent fallbacks to eager mode, introduce costly CPU-to-GPU synchronizations, and reduce optimization opportunities.
 GraphMend addresses this limitation by analyzing and transforming source code before execution.
-Built on the Jac compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python I/O functions.
+Built on the Jaseci compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python side effects.
 This design allows PyTorch’s compilation pipeline to capture larger, uninterrupted FX graphs without requiring manual refactoring by developers.
-Evaluation across eight Hugging Face models shows that GraphMend removes all fixable graph breaks due to dynamic control flow and Python I/O functions, driving the break count to 0 in 6 models and reducing it from 5 to 2 in another model.
+Evaluation across eight Hugging Face models shows that GraphMend removes graph breaks due to dynamic control flow and Python side effects, reducing the break count to 0 in 6 models and reducing it from 5 to 2 in another model.
 On NVIDIA RTX 3090 and A40 GPUs, GraphMend achieves up to 75% latency reductions and up to 8% higher end-to-end throughput.
 These results demonstrate that high-level code transformation is an effective complement to PyTorch’s dynamic JIT compilation pipeline, substantially improving both usability and performance.
-I Introduction
-PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility [1].
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
+1 Introduction
+PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility Paszke et al.
+(2019).
 However, its general-purpose design incurs efficiency overheads, leaving PyTorch workloads less optimized than handcrafted GPU implementations.
-To bridge this gap, PyTorch 2 [2] introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
-One of the central intermediate representations (IR) in this pipeline is the FX graph [3].
+To bridge this gap, PyTorch 2 Ansel et al.
+(2024) introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
+One of the central intermediate representations (IR) in this pipeline is the FX graph Reed et al.
+(2022).
 TorchDynamo intercepts and symbolically evaluates Python bytecode to create an FX graph.
-The FX graph is then passed as the IR to a backend compiler, by default TorchInductor [2], which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
-A key component in the PyTorch compilation pipeline is the capture of code into an FX graph through symbolic evaluation.
-It is important to avoid splitting the forward function into smaller graphs, since unsupported code triggers execution to fall back to eager mode, fragmenting the computation.
-For example, TorchDynamo often encounters difficulties when resolving conditional jumps, as control flow depending on dynamic values cannot be statically determined [4].
-In such cases like dynamic control flow or calls to certain Python built-ins, TorchDynamo cannot capture the affected code into the FX graph.
-Instead, it inserts a graph break, which splits the computation into multiple fragments [2].
-As a result, a single forward function may be compiled into several disjoint FX graphs rather than one unified graph.
-The compiled FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode [1].
-Each fallback to eager execution requires GPU kernels to synchronize with the CPU to fetch and schedule subsequent operations.
-These repeated host-device synchronizations introduce substantial overhead, since control is repeatedly transferred between CPU and GPU instead of executing long fused kernels entirely on the device.
-Moreover, graph breaks reduce optimization opportunities, as each FX graph is compiled in isolation, preventing cross-graph fusion and global optimization [5, 2].
+The FX graph is then passed to a backend compiler, by default TorchInductor Ansel et al.
+(2024), which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
+When tracing the FX graph, TorchDynamo often struggles with unsupported Python operations, which it handles by falling back to eager mode.
+These unsupported operations include dynamic control flow such as conditionals that depend on runtime values and certain Python side effects like I/O operations.
+In such cases, it inserts graph breaks, resulting in multiple disjoint FX graphs instead of a single unified one Zheng et al.
+(2023); Ansel et al.
+(2024).
+These disjoint FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode Paszke et al.
+(2019).
+Each fallback to eager mode forces GPU kernels to synchronize with the CPU to schedule subsequent operations, introducing significant overhead from repeated host-device communication.
+These synchronizations not only fragment execution but also reduce optimization opportunities, as each FX graph is compiled independently, preventing cross-graph fusion and global optimization Agrawal et al.
+(2019); Ansel et al.
+(2024).
 Avoiding graph breaks is essential for realizing the full performance potential of PyTorch 2.
-The current PyTorch compilation pipeline begins at the bytecode level and does not exploit source-level information.
-As a result, when high-level constructs such as conditional branches are lowered into jump instructions at the bytecode level, TorchDynamo often cannot statically resolve them, causing graph breaks.
-Our key insight is that compilers should not limit transformations to the bytecode level.
-Instead, more holistic compilation strategies that analyze and transform higher-level program representations closer to the source code can expose additional optimization opportunities.
-By incorporating source-level analysis and transformation, high-level code regions can be systematically rewritten into forms that enable more effective optimization and unified graph generation within the PyTorch 2 pipeline.
-Building on this insight, we introduce GraphMend, a high-level compiler technique for PyTorch 2 models that eliminates two common types of graph breaks those caused by dynamic control flow and Python I/O operations through program analysis and transformation.
-Unlike standard PyTorch 2, which captures execution at runtime to construct graphs for JIT compilation, GraphMend restructures Python code before execution.
-We implement GraphMend within the Jac framework [6, 7, 8], a language infrastructure that provides compiler and runtime support for Python programs.
-The Jac compiler constructs an abstract syntax tree (AST) to capture program structure and a control-flow graph (CFG) to represent execution paths, merging them into a unified IR [9].
-Multiple compiler passes operate on this IR to analyze and transform the program, and GraphMend extends this pipeline with additional passes that detect code patterns likely to cause graph breaks.
-Specifically, GraphMend introduces compiler passes that apply two key transformations: Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects, which systematically eliminate graph breaks.
-The transformed program is then compiled into standard Python bytecode and executed by the regular CPython interpreter, seamlessly integrating with the existing PyTorch 2 compilation pipeline.
-This design enables GraphMend to automatically restructure code before it reaches TorchDynamo.
-We evaluate GraphMend on a benchmark suite of 8 Hugging Face [10] models that trigger FX graph breaks.
-GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed.
-On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput.
-Improvements are larger when a graph break lands in the middle of a hot function where its disruption is highest, and smaller when the break is less disruptive; very small models also tend to see larger relative reductions in per forward pass latency because CPU to GPU handoff dominates a greater share of their runtime.
-Our contributions are, (1) We introduce a benchmark suite of eight Hugging Face models that naturally exhibit graph breaks, providing a realistic and reproducible testbed for evaluating PyTorch compilation.
-(2) We design and implement the GraphMend compilation technique, which automatically restructure high level code to eliminate common graph breaks.
-(3) We conduct comprehensive evaluations on this benchmark suite using real world models and hardware, demonstrating the effectiveness of our approach in reducing graph breaks and improving performance.
-II Background
-Optimization pipelines in PyTorch aim to capture execution graphs that can be globally optimized by compilers, unlike eager mode where each operation runs immediately and independently [1, 3].
-II-A Graph Capturing Techniques in PyTorch Before PyTorch 2
-Early approaches to graph capture were limited.
-torch.jit.trace [11] used record–replay at the C++ dispatcher level but ignored Python control flow and often produced incomplete or misleading graphs.
-torch.jit.script [12] analyzed Python ASTs but followed an “all-or-nothing” rule: if a single construct was unsupported, capture failed entirely.
-To address these limitations, PyTorch introduced the torch.fx [3] module, enabling symbolic execution of Python code to generate an FX graph, a higher-level IR.
-Although more flexible, symbolic_trace could still misrepresent certain operations.
-II-B FX Graph IR
-An FX graph represents operations as nodes connected by data dependencies [2, 3].
-Unlike eager execution, which evaluates operations sequentially, an FX graph exposes the computation as a data flow structure.
-This global view enables the compiler backend to apply optimizations such as operation fusion and kernel specialization.
-To illustrate, consider the simple PyTorch forward function in Figure 1(a).
-When captured by TorchDynamo, it is transformed into the FX graph shown in Figure 1(b), where each operation is represented as a node in the IR.
-1def forward(x, y):
-2 z = x*2 + y*2
-3 return torch.relu(z)
-(a) Forward function in PyTorch
-%x : torch.Tensor
-%y : torch.Tensor
-%mul_x : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%x, 2)
-%mul_y : torch.Tensor = torch.ops.aten.mul.Tensor_Scalar(%y, 2)
-%z : torch.Tensor = torch.ops.aten.add.Tensor(%mul_x, %mul_y)
-%out : torch.Tensor = torch.ops.aten.relu.default(%z)
-return (%out)
-(b) Corresponding FX graph IR
-Figure 1: A simple forward function (a) and its compiled FX graph IR (b).
-This FX graph serves as the foundation for subsequent compiler passes that optimize and lower the computation into efficient GPU kernels.
-II-C TorchDynamo in PyTorch 2
-PyTorch 2 [2] introduced TorchDynamo, Python bytecode-level tracer tightly integrated with CPython.
-It symbolically evaluates Python bytecode to extract PyTorch operations into FX graphs, while delegating unsupported constructs back to Python.
-This avoids incomplete graphs but may still fragment the computation by generating partial graphs by inserting graph-breaks.
+Some of these breaks can be eliminated by transforming Python code into forms that use PyTorch-supported operations, allowing the compiler to capture more of the program as a single, continuous FX graph.
+However, the current compilation pipeline analyzes code at the bytecode level, operating instruction by instruction without understanding higher-level program structures such as control flow, data dependencies, and side effects.
+Our key insight is that analyzing and transforming higher-level program representations closer to the source code, and integrating this with bytecode-level analysis, enables a more holistic compilation strategy that exposes richer optimization opportunities.
 1def f(x, y):
@@
 8 return torch.relu(z)
-(a) A PyTorch forward pass with data dependent control flow. The conditional branch on x.sum cannot be captured, causing a graph break.
+Figure 1: A PyTorch forward pass with data-dependent control flow.
+The conditional branch on x.sum cannot be captured, causing a graph break.
+(See also Figure 2 for the transformed version that eliminates the break.)
+Building on this insight, we introduce GraphMend, a high level compiler technique for PyTorch 2 that eliminates two major types of graph breaks, those arising from dynamic control flow and several Python side effects, through high-level program analysis and transformation.
+GraphMend extends the standard PyTorch 2 compilation pipeline with an AST-level transformation phase that restructures Python code before bytecode generation, improving traceability by eliminating graph breaks.
+We implement GraphMend within the Jaseci framework, built on Jac language Mars et al.
+(2023); Mars (2025); Jaseci Labs (2025), which is a super-set of Python.
+The Jac compiler constructs an abstract syntax tree (AST) and a control flow graph (CFG), merging them into a unified intermediate representation (UniiR) Dantanarayana et al.
+(2025).
+Multiple compiler passes analyze and transform this IR, and GraphMend extends this pipeline with additional passes that detect and restructure code patterns prone to graph breaks.
+Specifically, it introduces two transformations, Predicated Dynamic Control Flow and Graph Epilogue Deferred Side Effects, which systematically eliminate these breaks.
+We evaluate GraphMend on a benchmark suite of 8 Hugging Face Hugging Face (2025) models that trigger FX graph breaks.
+GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed.
+On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput.
+Our contributions are, (1) We introduce a benchmark suite of eight Hugging Face models that naturally exhibit graph breaks, providing a realistic and reproducible testbed for evaluating PyTorch compilation.
+(2) We design and implement the GraphMend compilation technique, which automatically restructure high level code to eliminate common graph breaks.
+(3) We conduct comprehensive evaluations on this benchmark suite using real world models and hardware, demonstrating the effectiveness of our approach in reducing graph breaks and improving performance.
 1def f(x, y):
@@
 10 return torch.relu(z)
-(b) Eliminating the graph break by rewriting data dependent control flow with torch.where, keeping the computation in a single FX graph.
-Figure 2: Comparison of control flow handling in torch.compile: (a) graph break due to Python control flow, and (b) fixed version using torch.where.
+Figure 2: Eliminating the graph break by rewriting data-dependent control flow with torch.where, keeping the computation in a single FX graph.
+Together with Figure 1, these illustrate how dataflow rewriting replaces dynamic Python branching with predicated tensor execution.
 Figure 3: Profiled traces of forward pass execution across CPU and GPU.
-(a) Forward pass execution trace of code with graph breaks in Figure 2(a).
-(b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2(b).
-III Motivating Example
-As discussed in § II, TorchDynamo attempts to capture a single FX graph for the forward function but inserts graph breaks whenever it encounters unsupported constructs, forcing execution to fall back to eager mode [2].
-III-A FX graph breaks
-Let us consider the modified program in Figure 2(a), which adds a conditional branch to the forward pass of Figure 1(a).
+(a) Forward pass execution trace of code with graph breaks in Figure 1.
+(b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2.
+(a) Existing graph-break in Phi-4-mini instruct model built by Microsoft.
+(b) Version of Phi-4-mini instruct model with the graph break fixed, without affecting program correctness.
+Figure 4: (a) Existing graph-breaks in a real model and (b) how it can be fixed to produce a single contiguous FX graph.
+2 Motivation
+FX graph capturing:
+TorchDynamo Ansel et al. (2024) captures graphs by symbolically interpreting Python bytecode, extracting PyTorch operations into FX graphs while falling back to eager mode for unsupported code.
+FX graph breaks.
+Let us consider the program in Figure 1, which has a conditional branch.
 When Dynamo tries to symbolically evaluate the function to capture FX graph, it cannot capture the conditional branch because the outcome depends on the runtime value of the tensor expression x.sum.
 Therefore, it inserts a graph break.
-As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph [2, 13].
-III-B Impact of graph breaks
-The backend compiler (TorchInductor by default) compiles each captured FX graph separately and lowers it to GPU executables such as fused Triton kernels [2, 14].
-Each compiled FX graph corresponds to a torch.compiled region.
-When graph breaks occur, a single Python function results in multiple compiled regions separated by segments of unsupported Python code [13].
+As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph Ansel et al.
+(2024); PyTorch Contributors (2025).
+Impact of graph breaks.
+When graph breaks occur, the function is divided into multiple FX graphs, each compiled separately into GPU kernels Ansel et al.
+(2024); Ghosh et al.
+(2025).
+Each FX graph forms a distinct torch.compiled region separated by segments of unsupported Python code PyTorch Contributors (2025).
+During execution, GPU runs the first compiled region.
+Control then returns to the Python interpreter to execute the unsupported code in eager mode and switches back to GPU for the next compiled region Ansel et al.
+(2024).
+These context switches introduce CPU–GPU synchronization overhead and reduce GPU utilization.
+Figure 3(a) illustrates this effect for the function in Figure 1.
+The trace shows two separate torch.compile regions divided by a device-to-host (D2H) memory transfer, where the GPU remains idle until the next CUDA graph launch.
+This fragmentation increases CPU scheduling overhead and prevents optimizations such as kernel fusion.
+When the break is removed and captured as a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA region with full GPU utilization and no intermediate D2H transfers.
+Kernel launches become streamlined, and synchronization overhead is eliminated.
+Real world model example with graph breaks.
+Phi-4-mini Instruct Microsoft (2025) is a lightweight open-source instruction-following model from Microsoft, available on Hugging Face as an efficient alternative to larger transformers.
+According to Hugging Face statistics Hugging Face (2025), the model recorded over 221k downloads in September 2025, reflecting strong community adoption.
+To examine the impact of graph breaks in a widely used model, we profiled Phi-4-mini Instruct using PyTorch 2 and identified 5 graph breaks caused by data-dependent control flow.
+Figure 4(a) illustrates this case, where conditional logic depending on runtime tensor values introduces a graph break.
+We rewrote the same control flow using PyTorch-compatible transformations with the torch.where operation to evaluate both branches within the FX graph while correctly handling side effects, as shown in Figure 4(b).
+After applying this modification, all graph breaks were eliminated.
+Profiling showed over a 50% latency reduction per forward pass in cold runs across multiple GPUs and more than a 5% improvement in steady-state runs compared to the version with graph breaks (further discussed in § 5), by reducing the impact of the graph breaks.
+This example demonstrates that targeted code transformations can effectively remove graph breaks and substantially improve performance.
+3 Understanding and Eliminating Common Graph Breaks
 1def fn(x):
@@
 4 return torch.sin(x)
-(a) Compiled function with a debug print of tensor statistics. Dynamo inserts a graph break at the print statement.
+Figure 5: Compiled function with a print statement. Dynamo inserts a graph break at the print statement. (See also Figure 6 for the transformed version that avoids the break.)
 1def fn(x):
@@
 6 return y
-(b) Fixing the graph break by registering print as a reorderable logging function. TorchDynamo can now safely reorder the call and keep the computation in a single FX graph.
-Figure 4: Fixing graph breaks due to Python I/O: (a) direct print causes a graph break, (b) reordering via variable assignment avoids the break.
-1 if x > 0:
-2 return 1
-3 else:
-4 return -1
-(a) Python function with if-else branch
-2 LOAD_CONST 1 (0)
-4 COMPARE_OP 4 (>)
-6 POP_JUMP_IF_FALSE 14
-8 LOAD_CONST 2 (1)
-10 RETURN_VALUE
-14 LOAD_CONST 3 (-1)
-16 RETURN_VALUE
-(b) Disassembled Python bytecode
-1 if x > 0:
-2 return 1
-3 return -1
-(c) Python function having identical bytecode to (a)
-Figure 5: Bytecode-level limitations of TorchDynamo.(a) (a) and (c) show two syntactically different functions with equivalent semantics. (b) Both compile to identical Python bytecode.
-During runtime, the first compiled region runs on the GPU, after which control returns to the Python interpreter to execute unsupported code in eager mode before switching back to the GPU for subsequent compiled regions.
-These context switches introduce overhead and lower GPU utilization per forward pass [2].
-This effect is illustrated in Figure 3(a), which shows the CPU–GPU profiling trace of the function in Figure 2(a).
-The trace reveals two separate torch.compile regions, corresponding to two FX graphs.
-The graph break is visible in the device-to-host (D2H) memcpy following the first FX graph, where control returns to the CPU to execute unsupported code, leaving the GPU idle until the second CUDA graph launch begins.
-This fragmentation both increases CPU-side scheduling overhead and prevents cross-graph optimizations such as kernel fusion.
-If the graph break is eliminated and the computation is captured in a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA graph region.
-In this case, the GPU remains fully utilized without falling back to the CPU, and no device-to-host memory transfers are observed between regions.
-As a result, kernel launches are streamlined and synchronization overhead is removed.
-Therefore, eliminating graph breaks is necessary.
-IV GraphMend Compiler Technique
+Figure 6: Fixing the graph break in Figure 5 by registering print as a reorderable logging function. The value to be printed is assigned to a buffer during tracing and flushed at the end.
 Figure 7: GraphMend compiler integration in the Jac pipeline.
@@
 GraphMend (bottom) analyzes entry points, detects graph breaks, and applies AST transformations (Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects) to eliminate breaks before execution.
-As discussed in § III, graph breaks in PyTorch 2 introduce significant overhead, undermining the performance gains expected in PyTorch 2.
-In this work, we focus on two common sources of graph breaks and introduce a compiler technique that eliminates them prior to execution.
-IV-A Classification of Graph Breaks
-In practice, most graph breaks arise from two common situations:
-  1. 1.
-    Data dependent operations [2, 13]: Constructs whose outcome depends on runtime values and hence, cannot be traced.
-  2. 2.
-    Unsupported operations [2, 13]: Several Python built in functions or certain functionalities that TorchDynamo does not support.
-In this section, we discuss these types of graph breaks.
-IV-A1 Data dependent operations
-Dynamo inserts graph breaks when it encounters control flow that depends on tensor values, such as if-statements or loops, as well as direct tensor data accesses (e.g., .item, .data_ptr).
-Among these cases, when Dynamo encounters dynamic control flow based on a tensor value as shown in Figure 2(a), we can fix the graph break by rewriting the code into operations supported by the GPU.
-IV-A2 Python builtin functions
-Some Python built-in functions such as printing, logging, or issuing warnings can also trigger graph breaks.
-Figure 4(a) shows an example where a debug print statement in the middle of the function causes a break.
-IV-B Fixing Common Graph Breaks
-Among these graph break reasons, there are two main reasons that are common and fixable which are (1) Data-dependent Control Flow and (2) Python I/O Operations.
-IV-B1 Data-dependent Control Flow
-Dynamo inserts graph breaks for constructs such as if-statements or loops where branching depend on tensor values.
-Returning to the example in Figure 2(a), the conditional introduces a graph break.
-In such cases,we can rewrite the function using tensor operations such as torch.where, which are supported by the GPU.
-The graph break fixed version using torch.where is shown in Figure 2(b) for the same function in Figure 2(a).
-This eliminates the graph break and allows Dynamo to capture a single FX graph.
-IV-B2 Python I/O Operations
-Python I/O functions such as Printing/logging/issuing warnings will result in a graph break.
-An example for this is shown in Figure 4(a) as there is a debug print statement in the middle of the function.
-If we can defer print statement to the end of the function, then we can avoid graph breaks as shown in Figure 4(b).
-These graph-fixing techniques are simple and straightforward: by applying source-level transformations, we can often avoid graph breaks.
-However, TorchDynamo traces at the bytecode level, which prevents it from leveraging such transformations and leads to unnecessary graph breaks.
-For example, consider Figure 5.
-The function in Figure 5(a) uses an explicit if–else block, while the function in Figure 5(c) omits the else and relies on a fall-through return.
-Although the two versions differ syntactically, both compile to the same bytecode shown in Figure 5(b).
-Because TorchDynamo operates purely on bytecode, it cannot distinguish between these semantically equivalent but syntactically distinct patterns.
-Similarly, logging operations introduce host I/O into the forward pass.
-Because it is unsafe to alter such effects directly at the bytecode level, TorchDynamo also treats these sites as graph breaks.
-In both cases, the limitation stems from relying solely on bytecode-level semantics captured just-in-time.
-To address this, we implement a compilation technique that analyzes and transforms programs at a higher-level IR.
-By operating on high-level IR, we preserve structural information lost in bytecode, enabling more sophisticated analyses and safe source-to-source rewrites that remove graph breaks before execution.
-This implementation of the technique is shown in Figure 6 which is elaborated in the rest of this section.
-IV-C Jac Compilation Pipeline
-To implement our compiler technique, we build on the Jac framework [6, 7, 9], which extends Python with additional language features through its runtime system.
-The Jac compiler accepts both Jac and Python code, parses it into an abstract syntax tree (AST), and progressively lowers it into a unified intermediate representation (UniiR).
-As illustrated in the top half of Figure 6, UniiR interleaves three structures: AST, control-flow graph (CFG), and symbol table.
-In this way, structural, semantic, and control-flow information are combined in a single representation.
-This integration enables cross-cutting analyses that are difficult to achieve when these structures are maintained separately.
-The CFG is constructed by linking AST nodes to their successors, embedding control-flow edges within the tree.
-This process is implemented in the CFGBuildPass.
-Our GraphMend pass takes UniiR as input and performs analyses to identify the two types of graph breaks, as illustrated in the bottom half of Figure 6.
-It then applies transformations on the AST to rewrite the code and eliminate these breaks.
-Afterward, UniiR is rebuilt and verified before being passed back to the Jac compiler pipeline.
-At that stage, the pipeline can either compile the program to bytecode for execution on the Python interpreter or generate a transformed Python source file.
+In this work, we focus on two common causes of graph breaks, outlining code transformations to eliminate them and explaining why bytecode-level fixes are challenging without higher-level IR context.
+Data dependent operations.
+Ansel et al.
+(2024); PyTorch Contributors (2025) In PyTorch 2, symbolic tracing builds FX graphs using FakeTensor objects, assuming control flow decisions can be inferred from static information such as tensor shapes or constants.
+When a condition depends on actual tensor data, (for example, if x.sum() > 0: in Figure 1), the compiler cannot safely evaluate it, so PyTorch inserts a graph break.
+We can fix them by expressing data-dependent branches using tensor operations such as torch.where or torch.cond, which are supported on GPU.
+Figure 1-2 shows the rewritten version of the same function, where the use of torch.where eliminates the graph break.
+Python Side-effects.
+Side effecting operations such as print or logging calls also cause graph breaks because they interact with the Python runtime and cannot be represented in a functional graph.
+Deferring these effects until after the graph executes can remove the break while preserving program behavior.
+In Figure 5 and 6, we show how moving a print statement to the end of the function allows the entire function to remain in one FX graph.
+Why automating these transformations at the bytecode level is challenging?
+Automating these transformations must occur before PyTorch’s bytecode analysis, but performing safe transformations at the bytecode level is challenging since much of the high-level semantic information is lost during compilation.
+When high-level constructs such as if/else or while loops are lowered to bytecode, they become low-level jump instructions like POP_JUMP_IF_FALSE and JUMP_BACKWARD.
+Since these opcodes are reused across different control-flow patterns, distinguishing loops from conditionals and identifying their bodies is non-trivial and often requires partial reverse engineering of the bytecode.
+Ensuring correctness during transformation further complicates this process.
+Moreover, performing data-flow analysis at the bytecode level is difficult since Python uses a stack-based rather than a register-based model.
+While variable operations themselves are explicit, intermediate expression values are pushed and popped from the operand stack without explicit naming, obscuring the flow of computed values without symbolic stack simulation.
+This simulation adds complexity and overhead to the analysis, whereas analyses on higher-level AST representations with explicit variable bindings and control flow are far simpler.
+Handling side effects at the bytecode level is also challenging.
+While delaying operations like built-in print or logger calls may be safe, Python’s dynamic nature allows users to redefine these functions.
+Determining whether a call refers to the original built-in or a user-defined override requires tracking symbol bindings, which is only possible with higher-level semantic context.
+A higher-level IR that retains program structure and data dependencies is therefore better suited for automatically applying these transformations.
+4 GraphMend Compiler Technique
+In this section we introduce our compiler technique GraphMend, which automates these code transformations using AST and CFG level analyses and AST-transformations, to eliminate graph breaks.
+The implementation of this technique is shown in Figure 7 which is elaborated in the rest of this section.
+4.1 Jac Compilation Pipeline
+To implement our compiler technique, we build on the Jaseci framework Mars et al.
+(2023); Mars (2025); Dantanarayana et al.
+(2025), a compiler stack designed for advanced analysis and transformation of Python programs.
+Jaseci compiler accepts both Jac and Python code, parses them into an AST, and progressively lowers the representation into UniiR.
+As shown in the top half of Figure 7, UniiR unifies the AST, CFG and symbol table into a single structure, combining syntactic, semantic, and control-flow information.
 Input: UniiR; TorchAttrTable (torch attributes $\mapsto$ {dynamic, static})
 Output: Tags: GraphBreak[DynCtrlFl], GraphBreak[logger/print]
-foreach Dynamo Entry Point $r$ in UniiR do
-  Initialize worklist with successors of $r$
-  while worklist not empty do
-     Pop node $n$ from worklist; push unseen successors of $n$
+1
+2foreach Dynamo Entry Point $r$ in UniiR do
+3   Initialize worklist with successors of $r$
+4 while worklist not empty do
+5      Pop node $n$ from worklist; push unseen successors of $n$
+6
     // 1) Dynamic control-flow via if-conditions
-     if $n$ is an IfStmt then
-        $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
-        $dyn\leftarrow\textbf{false}$
-        foreach $e\in E$ do
-           if $e$ has a torch attribute then
+7    if $n$ is an IfStmt then
+8       $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
+9       $dyn\leftarrow\textbf{false}$
+10       foreach $e\in E$ do
+11          if $e$ has a torch attribute then
              $a\leftarrow$ attribute of $e$
              // e.g., .sum
-              if TorchAttrTable[$a$] = dynamic then
-                 $dyn\leftarrow\textbf{true}$; break
-              end if
-           end if
-        end foreach
-       if not dyndyn then
-           if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
-              $dyn\leftarrow\textbf{true}$
-           end if
-        end if
-       if dyndyn then
-           Tag $n$ as GraphBreak[DynCtrlFl]
-        end if
-     end if
+12             if TorchAttrTable[$a$] = dynamic then
+13                $dyn\leftarrow\textbf{true}$; break
+14
+15
+16
+17       if not dyndyn then
+18          if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
+19             $dyn\leftarrow\textbf{true}$
+20
+21
+22       if dyndyn then
+23            Tag $n$ as GraphBreak[DynCtrlFl]
+24
+25
+26
     // 2) Logger / print side effects
-     if $n$ is a call to print or logger.* then
-        Tag $n$ as GraphBreak[logger/print]
-     end if
-  end while
-end foreach
+27    if $n$ is a call to print or logger.* then
+28         Tag $n$ as GraphBreak[logger/print]
+29
+30
+31
 Algorithm 1 Fixable Graph-Break Detection
-IV-D GraphMend Implementation
+4.2 GraphMend Implementation
 Our compiler technique consists of three main passes, illustrated in the bottom half of Figure 7.
-IV-D1 Dynamo Entry Point Analysis Pass
+Dynamo Entry Point Analysis Pass
 To identify potential TorchDynamo compilation entry points, this pass traverses the AST and analyzes each node.
 A node is tagged when a function or an nn.Module object wrapped or decorated with torch.compile is encountered.
-If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped. Otherwise, the tagged nodes serve as entry points for subsequent passes.
-IV-D2 Graph Break Type Analysis Pass
+If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped.
+Otherwise, the tagged nodes serve as entry points for subsequent passes.
+Graph Break Type Analysis Pass.
 We implement a compound analysis pass on UniiR to identify potential graph-break locations, following the procedure in Algorithm 1.
-Starting from each Dynamo entry point, the pass traverses the control-flow graph (CFG) to locate program points that may induce graph breaks.
+Starting from each Dynamo entry point, the pass traverses the CFG to locate program points that may induce graph breaks.
 For conditional branches, the pass inspects the subexpressions of the condition.
@@
 These annotations guide the subsequent transformation passes that eliminate graph breaks.
-IV-D3 AST Transformation Pass
-After identifying graph-break locations, the AST must be modified to eliminate them.
-As discussed earlier in this section, these fixes can be applied through source-level transformations.
-Following the same approach, the AST is modified according to the graph-break tags identified in the previous pass.
-Based on these tags, one of the following transformations is performed.
+AST Transformation Pass
+After identifying graph break locations, we modify the AST according to the graph break tags detected in the previous pass, after verifying the transformation is safe.
+Based on these tags, one of the following transformations is applied.
 Predicated Dynamic Control Flow Transformation
-To address data-dependent control-flow graph breaks, we first evaluate the branch condition and store it as a predicate.
-The control flow is then rewritten using torch.where, which enables predicated execution of both branches.
-This process is illustrated in Figure 6.
-In the AST Transformation Pass (Figure 6), the original branch if x.sum() > 5: ...
-else: ...
-is reduced by extracting the condition into a symbolic variable pred (R 1).
-Next, the two branch assignments are merged into a single dataflow expression using torch.where(pred, e1, e2).
-Finally, the if and else AST nodes are removed, and their child nodes are reconnected to the parent, producing a transformed tree in which both execution paths are explicitly represented.
-Graph-Epilogue Deferred Side Effect Transformation
-For print or logger calls, the message is first stored in a temporary variable.
-This variable is then passed to the outermost forward function and executed at the end of the function (epilogue) before returning, as shown in Figure 6.
-In the example ASTs in Figure 6, the original call print("Logger message") is rewritten as a deferred assignment deferred = "Logger message".
-This assignment is then moved to the epilogue block, where it is printed after graph execution (R 2).
-In this way, the logging effect is preserved while remaining safely outside the traced region.
-Note that any computation in the return statement must be moved before the epilogue, stored in a temporary variable, and then returned afterward to avoid introducing a graph break.
-IV-E Healing UniiR and Downstream Flow
-After the AST transformations, the UniiR is no longer consistent.
-Therefore, the symbol table and CFG must be rebuilt using the same passes that originally created them.
-As shown in Figure 6, once the GraphMend pass is complete, the SymTabBuildPass and CFGBuildPass are rerun to restore UniiR for downstream compilation.
+Before performing code transformation to address data-dependent control flow that causes graph breaks, we first analyze each branch body to identify which variables are modified and which functions are invoked in the if_true and if_false branches.
+If both branches produce the same output variables, their results can first be computed separately using two different temporary variables, and then combined with a torch.where expression that selects values based on the condition.
+If both branches call the same function, the function is moved outside the branches and executed only once, after its input arguments are merged using torch.where.
+The transformation proceeds by rewriting the AST to handle data-dependent control-flow graph breaks.
+The branch condition is first evaluated and stored as a predicate (pred).
+Each branch is then expressed as a dataflow computation producing temporary intermediate results, which are merged using torch.where(pred, e1, e2).
+After this reconciliation step, we handle common function calls ensuring that both their inputs and side-effect-free semantics are preserved.
+Finally, the original if and else nodes are removed, and their child nodes are reconnected to the parent, yielding a transformed AST in which both execution paths are represented explicitly as dataflow expressions.
+This process, illustrated in Figure 7, eliminates data-dependent control-flow constructs and enables predicated, differentiable execution compatible with graph compilation.
+Table 1: Models in benchmark suite
+Benchmark Model Name Graph break count Graph break reasons Total Operators
+biogpt Luo et al. (2022) 2 logger calls 408
+blenderbot-400M-distill Meta AI (2020) 3 logger calls 489
+flan-t5-large Chung et al. (2022) 3 logger calls 2130
+longformer-base-409 Beltagy et al. (2020) 5 logger calls, tensor.item() call 913
+moe-minicpm-x4-base babybirdprd (2025) 15 Dynamic shape operator 280
+Phi-4-mini-instruct Microsoft (2025) 5 Dynamic control flow 1541
+Qwen-Audio-Chat Chu et al. (2023) 2 Dynamic control flow 1647
+tiny-random-PegasusForCausalLM Face (2025) 2 logger calls 64
+Deferred Side Effect Transformation
+Before applying this transformation, we perform analysis to ensure that print and logger statements correspond to the built-in implementations rather than user-overloaded versions.
+This prevents interference with user-defined behavior or unintended semantic changes.
+Once verified, side-effecting calls such as print or logger are rewritten into deferred buffer assignments.
+Specifically, the value to be printed/ logged is stored in a dedicated runtime buffer together with a predicate flag indicating whether the branch containing the statement was executed.
+After the FX graph finishes execution, the buffer is flushed: each stored message is printed or logged only if its flag is true.
+This keeps side-effecting operations in the right order and with the same behavior, but safely outside the traced region.
 At the end of the Jac compiler pipeline, Python bytecode is generated for interpretation, enabling TorchDynamo to seamlessly operate on the transformed code.
-TABLE I: Models in benchmark suite
-Benchmark Model Name Graph break count Graph break reasons Total Operators
-biogpt [15] 2 logger calls 408
-blenderbot-400M-distill [16] 3 logger calls 489
-flan-t5-large [17] 3 logger calls 2130
-longformer-base-409 [18] 5 logger calls, tensor.item() call 913
-moe-minicpm-x4-base [19] 15 Dynamic shape operator 280
-Phi-4-mini-instruct [20] 5 Dynamic control flow 1541
-Qwen-Audio-Chat [21] 2 Dynamic control flow 1647
-tiny-random-PegasusForCausalLM [22] 2 logger calls 64
-V Evaluation
-We evaluate GraphMend on real-world models that exhibit graph breaks. Our goals are to quantify (i) how often GraphMend removes fixable breaks and (ii) the end-to-end improvement after those fixes.
-Experimental Setup
-We implement GraphMend on top of Jaseci [6, 7, 9], an open-source research infrastructure that provides both compiler and runtime control as a Python superset.
-Our experiments are conducted on NVIDIA RTX 3090 and A40 GPUs, using TorchInductor as the backend compiler.
-Benchmark Suite
-To construct our benchmark, we (1) randomly sample 65 Hugging Face [10] models spanning diverse architectures and task categories; (2) identify models that trigger FX graph breaks; and (3) retain the eight models with breaks for evaluation.
-Table I summarizes the resulting suite, including the number and causes of graph breaks and the total operator count.
-A similar methodology was used in prior works to construct benchmark suite [2].
-Profiling Methodology
-We profile the forward pass using PyTorch Profiler [23].
-For each model, we evaluate two configurations: original model and graph breaks fixed version.
-Each configuration is executed for 7 iterations: one cold-start iteration followed by six warm iterations.
-The profiler records CPU and CUDA activities and operator-level events; traces are exported and inspected with Chrome’s trace viewer.
-V-A GraphMend’s Ability to Fix Graph Breaks
-TABLE II: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
+5 Evaluation
+Table 2: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
 Benchmark Model Graph Breaks Fixed (%)
@@
 tiny-random-PegasusForCausalLM 2 100
-First, we check GraphMend’s ability to fix graph breaks across the benchmark suite.
-Table II shows that GraphMend fully eliminates graph breaks in 6 of the 8 models.
-It partially fixes longformer-base-4096 and leaves moe-minicpm-x4-base unchanged.
-By cause, GraphMend removes all breaks due to logging statements and data-dependent control flow.
-The remaining breaks: tensor.item() and dynamic-shape operators could not be fixed by code transformations.
-For the performance analysis that follows, we report results on the seven models whose break counts were fully eliminated or reduced, excluding moe-minicpm-x4-base.
-V-B Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
-We measure Relative Latency Improvement (%) as
-$100\times\frac{T_{\text{break}}-T_{\text{fixed}}}{T_{\text{break}}},$
-where $T_{\text{break}}$ is the per-forward-pass latency for the benchmarked model with graph breaks, and $T_{\text{fixed}}$ is the latency after applying GraphMend fixes.
-We report the relative improvement per forward pass for break-fixable models in the benchmark suite, using PyTorch Profiler data for both cold and warm (steady-state) runs.
-Cold Start Latency Improvement
-Figure 7(a) reports results on both RTX 3090 and A40 GPUs.
-Across models, we observe a relative latency reduction of about 30–75% for both GPUs, indicating substantial gains in cold runs.
-At cold start, each graph break incurs significant overhead because every CUDA graph requires a separate record and caching.
-After eliminating these breaks, only a single (or fewer) CUDA Graph capture is needed, resulting in fewer records and reduced capture time, which minimizes overhead.
-Consequently, we achieve larger relative latency improvements per forward pass in cold runs.
-A more detailed analysis is provided in  § V-D.
-Steady-State Latency Improvement
-Figure 7(b) shows steady-state improvements on RTX 3090 and A40 GPUs, ranging from about 2.5% to 25% across models.
-After the cold run, the captured graph is cached and replayed, removing the high capture overhead present initially.
-However, when graph breaks persist, execution still incurs GPU–CPU switches (as discussed in  § V-D), which add latency.
-Eliminating these breaks with GraphMend removes such switches, yielding additional improvements in relative latency per forward pass.
-Notably, tiny-random-PegasusForCausalLM shows over 25% improvement.
-Since this model is very small, with only 64 operators, the fixed costs from graph breaks (such as host–device synchronization, and extra graph launches) take up a large portion of the runtime.
-Removing these breaks therefore gives a huge speedup per forward pass.
 Figure 8: Latency improvements from GraphMend (a) Cold-start latency reductions (b) Steady-state latency reductions across benchmark models on RTX 3090 and A40 GPUs.
 Figure 9: Relative throughput improvement across benchmark suite
-V-C Relative Throughput Improvement by Fixing Graph Breaks
-We measure Relative Throughput Improvement (%) as
-$100\times\frac{Tp_{\text{fixed}}-Tp_{\text{break}}}{Tp_{\text{break}}},$
-where $Tp_{\text{break}}$ is the throughput (output tokens per second) for the model with graph breaks, and $Tp_{\text{fixed}}$ is the throughput after applying GraphMend fixes.
-We measure the relative throughput across the models in the benchmark suite that are fixable through GraphMend for both RTX 3090 and A40.
-Figure 7(b) shows a relative throughput improvement of about 5-8% across all models on both GPUs.
-This improvement follows from the relative latency reduction per forward pass observed in Section V-B: as each step becomes faster, the end-to-end tokens per second increase.
-The variation of improvement across models highlights an important point: the magnitude of improvement depends on the extent to which a graph break affects the forward path.
-When the impact of a break is small, the corresponding throughput gain is also limited.
-For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow.
-We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct from both GPUs.
-The difference arises because the break in Phi-4-mini-instruct occurs near the beginning of the forward function, whereas in Qwen-Audio-Chat it occurs in the middle, where its impact is higher.
-Thus, the location of a break determines its effect: fixing some breaks yields larger throughput gains, while fixing others has a smaller impact.
-tiny-random-PegasusForCausalLM achieves more than 25% lower forward-pass latency, but throughput gains are smaller.
-This is because throughput also depends on prefill, decode, CPU tasks (e.g., sampling, tokenization), and data transfers, which are unaffected.
-For such a small model, these fixed costs dominate, so removing the break speeds up GPU compute but has less impact on end-to-end tokens/s, especially when decode dominates generation.
-V-D Overhead Analysis of Graph Breaks
-We conducted an analysis to evaluate how fixing graph breaks with GraphMend reduces overhead. We analyze Qwen-Audio-Chat model running on an A40 GPU which has 2 graph breaks.
 (a) Profiler tracer for one iteration of warm runs for Qwen-Audio-Chat original model before fixing the graph break run in A40 GPU
@@
 Figure 10: (a) Profile traces.
-Figure 10: Profiler tracer of cold run of Qwen-Audio-Chat model run on A40 GPU.
 (a) Steady State - Original Model
@@
 Figure 11: CPU/GPU activity time traces for Qwen-Audio-Chat model run on A40 GPU.
-To ensure a consistent comparison in this section, all activity plots (Figures 11(a), 11(b), 11(c), and 11(d)) are drawn only for the time interval between the first and second CUDA graph executions in the original model; for the fixed model, we consider the identical time region.
+We evaluate GraphMend on real-world models that exhibit graph breaks. Our goals are to quantify (i) how often GraphMend removes fixable breaks and (ii) the end-to-end improvement after those fixes.
+Experimental Setup
+To implement our compiler technique, we build on Jaseci framework Mars et al.
+(2023); Mars (2025).
+Experiments are conducted on NVIDIA RTX 3090 and A40 GPUs using TorchInductor as the backend compiler.
+Benchmark Suite
+To construct our benchmark suite, we (1) randomly sample 65 Hugging Face Hugging Face (2025) models spanning diverse architectures and task categories; (2) test each model to identify whether it triggers FX graph breaks; and (3) retain all 8 models that exhibit graph breaks for evaluation.
+Table 1 summarizes the resulting suite, including the number and causes of graph breaks and the total operator count.
+A similar methodology was used in prior work Ansel et al.
+(2024).
+Profiling Methodology
+We profile the forward pass using PyTorch Profiler pyt (2021).
+We evaluate the original and graph-break–fixed versions of each model, running seven iterations with one cold start and six warm runs.
+The profiler records CPU, CUDA, and operator-level events, which are then analyzed using Chrome’s trace viewer.
+5.1 GraphMend’s Ability to Fix Graph Breaks
+First, we evaluate GraphMend’s effectiveness in fixing graph breaks across the benchmark suite.
+As shown in Table 2, GraphMend fully eliminates graph breaks in 6 models and resolves some graph breaks in 2 others.
+Graph breaks introduced by tensor.item() calls and dynamic-shape operators remain unfixed, as they cannot be resolved through our code transformations.
+For the performance analysis that follows, we report results on the 7 models where graph breaks were fully or partially eliminated, excluding moe-minicpm-x4-base.
+5.2 Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
+We measure the relative latency improvement as the percentage reduction in per-forward-pass latency after applying GraphMend fixes compared to the version with graph breaks.
+Results are reported for the models using PyTorch Profiler data from both cold and warm runs.
+Cold Start Latency Improvement
+Figure 8(a) shows results on RTX 3090 and A40 GPUs.
+Across models, we observe a 30–75% reduction in latency during cold runs.
+At cold start, each graph break introduces significant overhead as separate CUDA graphs must be recorded and cached.
+Eliminating these breaks reduces capture operations to a single (or fewer) CUDA graph, minimizing overhead and yielding larger per forward pass improvements.
+A detailed analysis appears in  § 5.4.
+Steady-State Latency Improvement
+Figure 8(b) shows steady state results on RTX 3090 and A40 GPUs, with 2.5–25% latency reductions across models.
+After the cold run, cached CUDA graphs eliminate capture overhead, but remaining graph breaks still cause GPU to CPU switches, adding latency.
+Removing these breaks with GraphMend eliminates such switches, improving forward pass performance.
+Notably, tiny-random-PegasusForCausalLM shows over 25% improvement.
+Since this model is very small (64 operators), the fixed costs from graph breaks (e.g.
+host–device synchronization, and extra graph launches) take up a large portion of the runtime.
+Removing these breaks therefore gives a huge speedup per forward pass.
+5.3 Relative Throughput Improvement by Fixing Graph Breaks
+We measure the relative throughput improvement as the percentage increase in output tokens per second after applying GraphMend fixes compared to the version with graph breaks.
+We measure relative throughput for all models on RTX 3090 and A40 GPUs.
+Figure 8(b) shows 5–8% improvements, consistent with the latency reductions in Section 5.2.
+As each forward pass becomes faster, overall tokens per second increase.
+The variation of improvement across models reflects how much a graph break affects the forward path: when the impact of a break is small, the corresponding throughput gain is also limited.
+For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow.
+We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct.
+The difference arises because the break in Phi-4-mini-instruct occurs early in the forward function, while in Qwen-Audio-Chat it occurs mid-function, where its impact is greater.
+Hence, a break’s location determines its effect, fixing some yields larger throughput gains than others.
+tiny-random-PegasusForCausalLM achieves more than 25% lower forward-pass latency, but throughput gains are smaller.
+This is because throughput also depends on prefill, decode, CPU tasks (e.g., sampling, tokenization), and data transfers, which are unaffected.
+For such a small model, these fixed costs dominate, so removing the break speeds up GPU compute but has less impact on end-to-end tokens/s, especially when decode dominates generation.
+5.4 Overhead Analysis of Graph Breaks
+We analyze how fixing graph breaks with GraphMend reduces overhead using the Qwen Audio Chat model on an A40 GPU, which contains two graph breaks.
+For a consistent comparison, all activity plots (Figures 11(a)11(d)) show the time interval between the first and second CUDA graph executions in the original model; the same interval is used for the fixed model.
 Steady-state run overhead analysis
 Figure 10(a) shows the profiler trace for the original model.
-Execution is split into three CUDA graphs because of two graph breaks, with idle GPU periods between the first and second FX graph executions.
-This is clearly visible in Figure 11(a), the GPU becomes idle while the CPU handles control flow in eager mode, including a device-to-host (D2H) memcpy triggered by dynamic control flow (graph break reason for this model).
+Execution is split into 3 CUDA graphs due to the 2 graph breaks, with idle GPU periods between the first and second FX graph executions.
+This is clearly visible in Figure 11(a), GPU becomes idle while CPU handles control flow in eager mode, including a D2H memcpy triggered by dynamic control flow.
 This CPU fallback introduces an overhead.
-After applying GraphMend, the rewritten code replaces the dynamic branch with a GPU-supported torch.where.
-The resulting trace (Figure 9(b)) shows one continuous CUDA graph.
+After applying GraphMend replacing torch.where, the resulting trace (Figure 10(b)) shows one continuous CUDA graph.
 As confirmed in Figure 11(c), GPU activity remains uninterrupted and CPU fallbacks disappear.
@@
 Cold runs introduce even larger overheads.
-In the profiler trace of original model (Figure 10), each graph break triggers a separate CUDA Graph recording and caching, leading to long idle GPU phases and heavy CPU involvement as shown in Figure 11(b).
-In this interval, we observe that GPU idle time dominates, while the CPU is heavily active, reflecting significant overhead.
-By contrast, the fixed model (Figure 11(d)) executes with a single capture, avoiding repeated records and reducing idle GPU time.
-Overall, GraphMend reduces overhead more significantly in cold runs than in steady state, explaining the substantial cold-start latency reductions per forward pass reported in Figure 7(a).
-V-E GPU Runtime Kernel Statistics Analysis
-We analyze how fixing graph breaks affects GPU kernel counts, ordering, and scheduling, and whether this enables new optimization opportunities.
-Using the Phi-4-mini-instruct model on an A40 GPU, the profiler reports 404 executed kernels for the original model and 393 for the fixed model with a reduction of 11.
-This drop indicates more fusion and fewer launches, as removing breaks allows the compiler to capture larger regions and emit fewer, larger kernels.
-Figure 12 illustrates this change.
-In the original model, 3 small kernels appear that are replaced in the fixed model by a single fused kernel.
-We also observe reordering: one kernel from the original run shifts upward in the fixed run.
-With one CUDA Graph, the runtime can reorder independent regions and pack launches more effectively, reducing stalls and improving locality.
-Graph breaks force separate CUDA Graph captures, blocking fusion across the break and restricting scheduling to smaller windows.
-Removing them allows larger portions of the forward pass to be captured together, enabling the compiler and runtime to fuse operations, remove redundant launches, and reorder independent work.
-As a result, kernel gaps shrink and device utilization improves.
+In the original model (Figure 11(b)), each graph break triggers a separate CUDA Graph recording and caching, causing long GPU idle phases and heavy CPU activity.
+GPU idle time dominates as the CPU handles repeated captures.
+In contrast, the fixed model (Figure 11(d)) performs a single capture, minimizing GPU idle time and CPU work.
+Consequently, GraphMend yields larger overhead reductions in cold runs, explaining the substantial cold-start latency gains per forward pass shown in Figure 8(a).
+5.5 GPU Runtime Kernel Statistics Analysis
+We examine how fixing graph breaks affects GPU kernel execution and optimization.
+On Phi-4-mini-instruct model (A40 GPU), the profiler reports 404 kernels for the original model and 393 for the fixed one; a reduction reflecting increased fusion and fewer launches as larger regions are captured.
+As shown in Figure 12, three small kernels in the original model merge into a single fused kernel after fixing.
+Kernel reordering also appears, as the unified CUDA Graph enables better packing and scheduling of independent regions, reducing stalls and improving locality.
+Graph breaks previously forced separate captures, preventing cross-break fusion and limiting scheduling windows.
+Their removal allows broader fusion, redundant launch elimination, and more flexible reordering; reducing kernel gaps and boosting device utilization.
 In summary, fixing graph breaks with GraphMend not only reduces launch overhead but also exposes new optimization opportunities, underscoring the importance of eliminating such breaks.
 Figure 12: Kernel Fusion and Reordering Visualization for Phi-4-mini-instruct model.
-VI Related Work
+6 Related Work
+Table 3: Comparison of GraphMend with related transformation frameworks.
+Feature AutoGraph MagPy TensorSSA GraphMend
+Moldovan et al. (2019) Zhang et al. (2024) Ma et al. (2024)
+[…]
+Fully compatible with PyTorch 2 pipeline $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
+Handling side effects compatible with PyTorch $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
+[…]
+[…]
 PyTorch has long sought to balance performance with Python’s flexibility.
-TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs [12, 11].
-PyTorch 2 introduced TorchDynamo for bytecode-level capture and TorchInductor for backend compilation [2, 1], yet graph breaks remain common when encountering data-dependent branches or Python I/O.
-Our work reduces these breaks by rewriting source code before execution, allowing more continuous FX graphs.
-General compiler infrastructures such as XLA [24], TVM [25], and MLIR [26] provide IR-based optimization, while Glow [27], TASO [28], and ONNX Runtime [29] extend portability across hardware and frameworks.
-These systems assume large graphs are already available, whereas we restructure PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
-Dynamic models motivate systems like GRAPE [30] and Triton [31, 32], which manage runtime variability through efficient kernel generation.
-PyTorch’s compiler stack still falls back to eager execution in such cases, while our method removes common break sources by rewriting branches into tensorized forms that remain graph-compatible.
-Program transformations more broadly, as seen in Halide [33], Lantern [34], and DLVM [30], show the potential of semantics-preserving rewrites for optimization.
-Our approach follows this tradition but targets PyTorch specifically, introducing lightweight transformations that enable TorchDynamo to capture larger uninterrupted graphs.
-VII Conclusion
-We presented GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution.
-By restructuring code constructs such as dynamic control flow and Python I/O operations, GraphMend reduces fragmentation in the compilation pipeline and enables larger unified graphs to be captured by TorchDynamo.
-Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable breaks and consistently improves runtime efficiency, with substantial latency and throughput gains on modern GPUs.
+TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs PyTorch-TorchScript Team (2025); PyTorch Team (2025).
+PyTorch 2 introduced TorchDynamo for bytecode-level graph capture Ansel et al.
+(2024); Paszke et al.
+(2019), yet data-dependent branches and Python side effects still cause frequent graph breaks.
+GraphMend mitigates these by transforming the Python AST before execution.
+While AutoGraph Moldovan et al.
+(2019) generates computation graphs in TensorFlow through source-level transformations, its approach does not generalize to PyTorch 2, which captures graphs at the bytecode level.
+In contrast, GraphMend performs lightweight AST-level transformations integrated directly into the PyTorch 2 compilation pipeline, improving symbolic evaluation and graph completeness.
+MagPy Zhang et al.
+(2024) takes a different route by replacing TorchDynamo and reconstructing graphs from runtime traces, whereas GraphMend preserves the standard compilation flow with minimal overhead.
+TensorSSA Ma et al.
+(2024) introduces a custom SSA-based IR to manage tensor mutations and control flow but adds significant infrastructure and omits Python side effects handling.
+GraphMend instead applies predicate-based SSA-inspired rewrites directly on the Python AST, avoiding extra IR layers while staying fully compatible with PyTorch’s ecosystem.
+More broadly, GraphMend’s source-level rewriting follows the tradition of semantics-preserving transformations in program optimization, as exemplified by Halide Ragan-Kelley et al.
+(2013), Lantern Wang et al.
+(2018), and DLVM Zhao et al.
+(2017).
+Like these systems, GraphMend leverages structural code analysis for performance gains but is uniquely tailored to PyTorch, enabling TorchDynamo to capture larger, uninterrupted graphs.
+General compiler infrastructures such as XLA Community (2023), TVM Chen et al.
+(2018), and MLIR Lattner et al.
+(2020) provide IR-based optimization, while Glow Rotem et al.
+(2018), TASO Jia et al.
+(2019), and ONNX Runtime Microsoft (2018) extend portability across hardware and frameworks.
+These systems assume large graphs are already available, whereas GraphMend restructures PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
+7 Conclusion
+We present GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution.
+By restructuring code constructs such as dynamic control flow and Python side effects (e.g., I/O operations), GraphMend reduces fragmentation in the compilation pipeline and enables TorchDynamo to capture larger, unified graphs.
+Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable graph breaks and consistently improves runtime efficiency, achieving substantial reductions in latency and increases in throughput on modern GPUs.